OpenStudy (anonymous):

Solve the following equation 10/3 x= -50

OpenStudy (anonymous):

(10/3)(x) =-50 (10/3)(x)(3/10)=(-50)(3/10) <--- to make x on LHS and numbers on RHS x=-15

OpenStudy (anonymous):